OpenStudy (anonymous):

The perimeter of a rectangle is four more than the area of the rectangle. Find the area of the rectangle, if its width is 1 cm. What is the surface area of a cylinder that has a volume of 785.4 in^3 and a radius of 5 in?

OpenStudy (anonymous):

Let x = height Area = x*1 = x Perimeter = 2x + 2*1 = 2x + 2 Perimeter = Area + 4 so 2x + 2 = x + 4 2x - x = 4 - 2 x = 2 Therefore, height = 2cm Therefore, Area = 2*1 = 2cm^2

OpenStudy (anonymous):

the choices of the first one are : 0.5 cm^3, 2cm^3, 8cm^3, 1cm^3

OpenStudy (anonymous):

area(circular side) = pi*5^2 = 25pi V = area(circular side) * length therefore, length = V/area(circular side) = 785.4/25pi = 10 S.A. (Surface Area) = 2*area(circular side) + area(length side) = 2*25pi + length*circumference = 50pi + 10*2pi*5 = 50pi + 100pi = 150pi in^2 = 471.2 in^2

OpenStudy (shrutipande9):

ahh my bad pheonix is right....

OpenStudy (anonymous):

is the first one 2cm^3?

OpenStudy (shrutipande9):

no...if its area..then it is sq cm or cm^2

OpenStudy (anonymous):

ok. Thanks